THE SUM OF THE AGES OF SON AND FATHER IS 56 YERAS.AFTER 4 YEARS THE AGE OF THE FATHER WILL BE THREE TIMES THAT OF THE SON.FIND THE AGE OF THE FATHER?

Answer by lwsshak3(11628)

THE SUM OF THE AGES OF SON AND FATHER IS 56 YERAS.AFTER 4 YEARS THE AGE OF THE FATHER WILL BE THREE TIMES THAT OF THE SON.FIND THE AGE OF THE FATHER?
**
let x= age of father
56-x=age of son
after 4 yrs
x+4=age of father
(56-x)+4=(60-x) age of son
x+4=3(60-x)
x+4=180-3x
4x=176
x=44
age of father: 44